OpenStudy (anonymous):

Consider the following pair of equations: −x − y = −5 y = x + 1 If the two equations are graphed, at what point do the lines representing the two equations intersect? (−2, 3) (3, −2) (2, 3) (3, 2)

OpenStudy (anonymous):

WHAT DO U THINK

OpenStudy (anonymous):

b

OpenStudy (kidrah69):

\[-x-(x=1)=-5\]

OpenStudy (anonymous):

I have no idea

OpenStudy (kidrah69):

distribute the negative 1st

OpenStudy (kidrah69):

\[-(x+1)\]

OpenStudy (anonymous):

- + -1

OpenStudy (kidrah69):

-x-1

OpenStudy (anonymous):

-x + -1 **

OpenStudy (kidrah69):

\[-x-x-1=-5\]

OpenStudy (anonymous):

-x -1 = -5

OpenStudy (kidrah69):

whats -x-x?

OpenStudy (anonymous):

idk D:

OpenStudy (kidrah69):

ok you are technically subtracting from -1 therefore you get -2 now you have \[-2x-1=-5\]

OpenStudy (kidrah69):

do you know what to do next?

OpenStudy (anonymous):

-5+-1?

OpenStudy (kidrah69):

no never put +-1 it makes no sense but you have the right idea what is -5+1?

OpenStudy (anonymous):

-4

OpenStudy (kidrah69):

good! now what is : \[-2x/-4\]

OpenStudy (anonymous):

2

OpenStudy (anonymous):

x = 2

OpenStudy (kidrah69):

yep i think you can determine what your answer will be from that :)

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Refer to the attached Mathematica plot.